Calculate the energy required to raise the temperature of 11.4 grams of gaseous nitrogen from 23.0 °C to 36.4 °C. Answer:
Answer #1

q= mass * specific heat capacity * CHANGE IN TEMPERTATURE

GIVEN

MASS= 11.4 GRAMS= 11.4*10^-3 Kg

SPECIFIC HEAT CAPACITY = 1.039 KJ/kg.K

CHANGE IN TEMPERATURE = (36.4-23.0) =13.4K

q= 11.4*10^-3*1.039*13.4

q= 0.15871KJ

q= 158.71 J